OpenStudy (anonymous):

A car transporter can carry 12 cars How many transporters are need for 67 cars ?

OpenStudy (jhannybean):

\[\sf \frac{1~transporter}{12~cars} = \frac{x~transporters}{67~cars}\]

OpenStudy (texaschic101):

1/12 = x/67 --- 1 transporter to 12 cars = x transporters to 67 cars now we cross multiply 12x = 67 -- now divide both sides by 12 x = 67/12 x = 5 7/12.....um....it can carry 5 cars...not sure about the 7/12...thats not even a whole car

OpenStudy (anonymous):

So the answer is 5?